# Quantum Speedups in Regret Analysis of Infinite Horizon Average-Reward Markov Decision Processes

### v2

@@ VI Conclusions and Future Work
Appendix A Convexity of Optimization Problem
In order to solve the optimistic optimization problem in Eq. (16) - (20) for epoch $e$, we adopt the approach proposed in [9, 37]. The key is that the constraint in (19) seems non-convex. However, that can be made convex with an inequality as
@@ Appendix A Convexity of Optimization Problem
Appendix B A brief overview of Quantum Amplitude Amplification
@@
    Quantum Search: In the quantum algorithm for unstructured search, known as Grover’s algorithm [38], the technique of amplitude amplification is employed to enhance the amplitude of the target state and decrease the amplitude of the non-target states. As a result, this technique provides a quadratic improvement over classical search algorithms.
@@
    Quantum Optimization: Quantum amplification can be used in quantum optimization algorithms, such as Quantum Approximate Optimization Algorithm (QAOA) [39], to amplify the amplitude of the optimal solution and reduce the amplitude of sub-optimal solutions. This can lead to an exponential speedup in solving combinatorial optimization problems.
@@
    Quantum Simulation: Quantum amplification has also been used in quantum simulation algorithms, such as Quantum Phase Estimation (QPE) [40], to amplify the amplitude of the eigenstate of interest and reduce the amplitude of other states. This can lead to an efficient simulation of quantum systems, which is an intractable problem in the classical domain.
@@
    Quantum Machine Learning: The utilization of quantum amplification is also evident in quantum machine learning algorithms, including Quantum Support Vector Machines (QSVM) [41]. The primary objective is to amplify the amplitude of support vectors while decreasing the amplitude of non-support vectors, which may result in an exponential improvement in resolving particular classification problems.
In the domain of quantum Monte Carlo methods, implementing quantum amplitude amplification can enhance the algorithm’s efficiency by decreasing the amount of samples needed to compute the integral or solve the optimization problem. By amplifying the amplitude of the desired state, a substantial reduction in the number of required samples can be achieved while still maintaining a certain level of accuracy, leading to significant improvements in efficiency when compared to classical methods. This technique has been particularly useful in [14] for achieving efficient convergence rates in quantum Monte Carlo simulations.

### v3

@@ 6 Conclusions and Future Work
A Convexity of Optimization Problem
In order to solve the optimistic optimization problem in Eq. (15) - (19) for epoch $e$, we adopt the approach proposed in (?, ?). The key is that the constraint in (18) seems non-convex. However, that can be made convex with an inequality as
@@ A Convexity of Optimization Problem
B A brief overview of Quantum Amplitude Amplification
@@
    Quantum Search: In the quantum algorithm for unstructured search, known as Grover’s algorithm (?), the technique of amplitude amplification is employed to enhance the amplitude of the target state and decrease the amplitude of the non-target states. As a result, this technique provides a quadratic improvement over classical search algorithms.
@@
    Quantum Optimization: Quantum amplification can be used in quantum optimization algorithms, such as Quantum Approximate Optimization Algorithm (QAOA) (?), to amplify the amplitude of the optimal solution and reduce the amplitude of sub-optimal solutions. This can lead to an exponential speedup in solving combinatorial optimization problems.
@@
    Quantum Simulation: Quantum amplification has also been used in quantum simulation algorithms, such as Quantum Phase Estimation (QPE) (?), to amplify the amplitude of the eigenstate of interest and reduce the amplitude of other states. This can lead to an efficient simulation of quantum systems, which is an intractable problem in the classical domain.
@@
    Quantum Machine Learning: The utilization of quantum amplification is also evident in quantum machine learning algorithms, including Quantum Support Vector Machines (QSVM) (?). The primary objective is to amplify the amplitude of support vectors while decreasing the amplitude of non-support vectors, which may result in an exponential improvement in resolving particular classification problems.
In the domain of quantum Monte Carlo methods, implementing quantum amplitude amplification can enhance the algorithm’s efficiency by decreasing the amount of samples needed to compute the integral or solve the optimization problem. By amplifying the amplitude of the desired state, a substantial reduction in the number of required samples can be achieved while still maintaining a certain level of accuracy, leading to significant improvements in efficiency when compared to classical methods. This technique has been particularly useful in (?) for achieving efficient convergence rates in quantum Monte Carlo simulations.